Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probo

Expert replies
by maihuna » Tue Mar 15, 2011 12:12 pm
In a certain lottery, the probability that a number between 12
and 20, inclusive, is drawn is 1
6 . If the probability that a
number 12 or larger is drawn is 2
3 , what is the probability
that a number less than or equal to 20 is drawn?
(A) 1
18
(B) 1
6
(C) 1
3
(D) 1
2
(E) 5
6
Charged up again to beat the beast :)
Join the discussion
Source: — Problem Solving |

by force5 » Tue Mar 15, 2011 1:55 pm
please check the question.. its not making sense.... some symbols are missing somewhere... its that 1/6 and 2/3 or what??
Join the discussion

by Stuart@KaplanGMAT » Tue Mar 15, 2011 2:09 pm
maihuna wrote:In a certain lottery, the probability that a number between 12 and 20, inclusive, is drawn is 1 in 6 . If the probability that a number 12 or larger is drawn is 2 in 3, what is the probability that a number less than or equal to 20 is drawn?

(A) 1/18
(B) 1/6
(C) 1/3
(D) 1/2
(E) 5/6
I fixed the question - please always preview your posts to make sure that they're readable!

This is a simple question to solve once you've broken it down, so let's get to breaking!

If the probability of 12 or larger is 2/3 and the probability of 12-20 is 1/6, then we can set up and solve the following equation:

Prob(12+) = Prob(12-20) + Prob(21+)
2/3 = 1/6 + Prob(21+)
2/3 - 1/6 = Prob(21+)
4/6 - 1/6 = Prob(21+)
3/6 = 1/2 = Prob(21+)

We can now answer the question by setting up and solving one more equation:

Prob(want) = 1 - Prob(don't want)
Prob(less than or equal to 20) = 1 - Prob(greater than 20)
Prob(less than or equal to 20) = 1 - 1/2 = 1/2

choose (D)!
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion